# Equilibrium in Style: A Modeling Framework on the Cash Flow and the Life Cycle of a Consumer Store

### v1

@@ 8.2 Challenges in the empirical design
Several challenges are encountered during the empirical testing of our model. First, in Section 7, we use the category (or the sub-industry) to filter the competitors, implying that the stores in the same category own a similar style. However, this may result in some problems. For example, the stores of Miishien and Hangzhou soup dumplings stores belong to the same category of “simple food”, but it is not reliable that they are indistinguishable styles for consumers. Therefore, we may have overestimated the number of competitors for each store in our empirical research, thus underestimating the potential customers and overestimating the initial conversion rate. In empirical results, the initial conversion rate of store B from the restaurant industry is only slightly lower than that of store A from the retail industry may because of overestimation of store B competitors and underestimate the potential consumers.
@@
Finally, in real life scenarios, when a store’s cash flow dips below the break-even point, owners may not promptly opt for closure of their store. Figure 10 depicting store C highlights the owner’s persistent efforts to bolster cash flow, which ultimately proved unsuccessful. In practice, owners typically dedicate an extended period to enhancing cash flow, often through methods like discounted promotions. This period is usually three months in China, because of quarterly rent obligations.
@@ 8.3 Comments on the model settings
There are also some settings of our theoretical model that are worth emphasizing and discussing in detail. First, in 3.3, we show that even stores that open with the same initial monetary investment in the same time may choose different styles, which is consistent with real life. However, we assume a probability ordering rule to simplify in order to derive the Nash Equilibrium. Even if it is not assumed, we suspect that there should be an equilibrium in this style market, possibly a mixed equilibrium. Second, one might argue that our assumption population structure does not change over time, but in fact the assumption of preference shifting implies the explanation. What is more likely to happen in reality is that both population structure and consumers’ preferences change, but we can always encapsulate common preference groups that are stable over time into more specific types (the most specific case, each individual is a type). The only thing we need to focus on is how the preferences of the specific types change. Third, the preference shifting of consumers is depicted by the classic Cobb-Douglas utility function in Assumption 1. Nevertheless, the theory of behavioral economics has become more and more popular and is shown to be more realistic; see Barberis and Thaler (2003). E.g., the cumulative prospect theory (Tversky and Kahneman (1992)) and related S-shaped utility functions with new variants (Liang and Liu (2020), Liang et al. (2024)) could be explored to calibrate the customers’ preferences. Lastly, the Nash Equilibrium derived in this paper relies on the existence of a type of dominant consumers with a high proportion. If this is not the case, the store’s decisions with respect to consumers’ behavior may be discontinuous. There are counterexamples that prove that the equilibrium in Theorem 4.1 is not guaranteed under the existing conditions. The main reason for this is that it is hard to ensure that optimal purchase probability density for each type of consumer is simultaneously normalized.
@@ 9 Concluding remarks
The structural model we present offers insight into the cash flow generation processes at the micro level, providing explanations applicable to practical scenarios. Notably, while most stores exhibit a standard rainbow-shaped cash flow pattern, anomalies such as an “M” shape can be accounted for by considering diverse consumer preferences in our model. Besides, we find that a faster visibility broadening speed not only accelerates ramp-up periods but also amplifies cash flow peaks (Figure 10). This suggest that store owners should be encouraged to increase promotional activities during store openings. Furthermore, the decrease coefficient $\nu$ emerges as a pivotal parameter shaping a store’s life cycle dynamics, wherein regular menu updates or periodic location changes can mitigate the shifting effect of consumer preference. China’s bubble tea stores, such as NAIXUE listed in the Hong Kong exchange market, update their products per month. In this regard, policymakers may consider supporting innovative products to foster sustained market competitiveness. To conclude, this paper provides an answer to the open question of why the typical store usually has such a short life cycle.
@@
Acknowledgements. The authors thank Zhuo Chen, Zherui Fan, Chengxin Gong, Yangbo He, Zhiyi Song, Ningxin Zhang, Zhen Zhou, Elizabeth Zhu and members of the research group on financial mathematics and risk management at CUHK-Shenzhen for their helpful discussion and important feedback. Y. Liu acknowledges financial support from the research startup fund at The Chinese University of Hong Kong, Shenzhen (Grant No. UDF01003336).
@@ Appendix A Details in definition
A.1 Definition of purchase probability density when $\text{dim}(\mathbb{Z}_{t})<p+q.$
Denote the Lebesgue measure (or Hausdorff measure) on $\mathbb{Z}_{t}$ as $\sigma_{\mathbb{Z}_{t}},$ The consumers’ purchase probability measure on $\mathbb{Z}_{t}$ is $\mathbb{P}_{jt}.$ Then we define
@@
and the integral operation with respect to it is defined as
@@ Appendix B Proofs
From Proposition 41 we know that
@@
where $e_{\min}\left(H^{-1}_{L_{m}}\left(\boldsymbol{z},\theta\right)\right)$ represents smallest eigenvalue of $H^{-1}_{L_{m}}\left(\boldsymbol{z},\theta\right),$ see Strang (2012). It can be seen that
@@
Still we assume that $t=0$ without loss of generality. Given $K_{10},$ we show first that this is true for $B_{t}(\cdot)$ in Proposition 3.5. From the proof in Proposition 3.5, we know that

### v2

@@ 8.2 Challenges in the empirical design
Several challenges are encountered during the empirical testing of our model. First, in Section 7, we use the category (or the sub-industry) to filter the competitors, implying that the stores in the same category own a similar style. However, this may result in some problems. For example, the stores of Miishien and Hangzhou soup dumpling stores belong to the same category of “simple food”, but it is not reliable that they are indistinguishable styles for consumers. Therefore, we may have overestimated the number of competitors for each store in our empirical research, thus underestimating the potential customers and overestimating the initial conversion rate. In empirical results, the initial conversion rate of store B from the restaurant industry is only slightly lower than that of store A from the retail industry may be because of the overestimation of store B’s competitors and underestimation of the potential consumers.
@@
Finally, in real-life scenarios, when a store’s cash flow dips below the break-even point, owners may not promptly opt for the closure of their store. Figure 10 depicting store C highlights the owner’s persistent efforts to bolster cash flow, which ultimately proved unsuccessful. In practice, owners typically dedicate an extended period to enhancing cash flow, often through methods like discounted promotions. This period is usually three months in China, because of quarterly rent obligations.
@@ 8.3 Comments on the model settings
There are also some settings of our theoretical model that are worth emphasizing and discussing in detail. First, in Proposition 3.3, we show that even stores that open with the same initial monetary investment at the same time may choose different styles, which is consistent with real life. However, we assume a probability ordering rule to simplify to derive the Nash Equilibrium. Even if it is not assumed, we suspect that there should be an equilibrium in this style market, possibly a mixed equilibrium. Second, one might argue that our assumption of population structure does not change over time, but in fact, the assumption of preference shifting implies the explanation. What is more likely to happen in reality is that both population structure and consumers’ preferences change, but we can always encapsulate common preference groups that are stable over time into more specific types (in the most specific case, each individual is a type). The only thing we need to focus on is how the preferences of the specific types change. Third, the preference shifting of consumers is depicted by the classic Cobb-Douglas utility function in Assumption 1. Nevertheless, the theory of behavioral economics has become more and more popular and is shown to be more realistic; see Barberis and Thaler (2003). E.g., the cumulative prospect theory (Tversky and Kahneman (1992)) and related S-shaped utility functions with new variants (Liang and Liu (2020), Liang et al. (2024)) could be explored to calibrate the customers’ preferences. Lastly, the Nash Equilibrium derived in this paper relies on the existence of a type of dominant consumers with a high proportion. If this is not the case, the store’s decisions with respect to consumers’ behavior may be discontinuous. Some counterexamples prove that the equilibrium in Theorem 4.1 is not guaranteed under the existing conditions. The main reason for this is that it is hard to ensure that the optimal purchase probability density for each type of consumer is simultaneously normalized.
@@ 9 Concluding remarks
The structural model we present offers insight into the cash flow generation processes at the micro level, providing explanations applicable to practical scenarios. Notably, while most stores exhibit a standard rainbow-shaped cash flow pattern, anomalies such as an “M” shape can be accounted for by considering diverse consumer preferences in our model. Besides, we find that a faster visibility broadening speed not only accelerates ramp-up periods but also amplifies cash flow peaks (Figure 10). This suggests that store owners should be encouraged to increase promotional activities during store openings. Furthermore, the decreasing coefficient $\nu$ emerges as a pivotal parameter shaping a store’s life cycle dynamics, wherein regular menu updates or periodic location changes can mitigate the shifting effect of consumer preference. China’s bubble tea stores, such as NAIXUE listed in the Hong Kong exchange market, update their products per month. In this regard, policymakers may consider supporting innovative products to foster sustained market competitiveness. To conclude, this paper provides an answer to the open question of why the typical store usually has such a short life cycle.
@@
Acknowledgements. The authors thank Zhuo Chen, Zherui Fan, Chengxin Gong, Yangbo He, Zhiyi Song, Ningxin Zhang, Zhen Zhou, Elizabeth Zhu, and members of the research group on financial mathematics and risk management at CUHK-Shenzhen for their helpful discussion and important feedback. Y. Liu acknowledges financial support from the research startup fund at The Chinese University of Hong Kong, Shenzhen (Grant No. UDF01003336).
@@ Appendix A Details in definition
A.1 Definition of purchase probability density.
As long as $\text{dim}(\mathbb{Z}_{t})>0,$ we can define the purchase probability density. Denote the Lebesgue measure (or Hausdorff measure) on $\mathbb{Z}_{t}$ as $\sigma_{\mathbb{Z}_{t}},$ The type $j$th consumers’ purchase probability measure on $\mathbb{Z}_{t}$ is $\mathbb{P}_{jt},$ which satisfies that $\mathbb{P}_{jt}\ll\sigma_{\mathbb{Z}_{t}}.$ Then we define
@@
and the integrals with respect to it is defined as
@@ Appendix B Proofs
From Proposition 41, we know that
@@
where $e_{\min}\left(H^{-1}_{L_{m}}\left(\boldsymbol{z},\theta\right)\right)$ represents the smallest eigenvalue of $H^{-1}_{L_{m}}\left(\boldsymbol{z},\theta\right)$; see Strang (2012). It can be seen that
@@
Still, we assume that $t=0$ without loss of generality. Given $K_{10},$ we show first that this is true for $B_{t}(\cdot)$ in Proposition 3.5. From the proof in Proposition 3.5, we know that